OpenStudy (anonymous):

Medal! Input (x) 2 4 6 8 Output (y) 1 2 3 4 Part A: Do the data in this table represent a function? Justify your answer. Part B: Compare the data in the table with the relation f(x) = 3x – 10. Which relation has a greater value when x = 8? Part C: Using the relation in Part B, what is the value of x if f(x) = 80?

OpenStudy (error1603):

PART A) every x values goes t0 ONLY one y value A function is when x values does not repeat , if there is same x value more than onc than relation is a not function PART B) substitute x value and then solve for f(x) in the table when x = 8 y =what ? for example when x=6 y =3 PART C) substitute 80 for f(x) and then solve for x

OpenStudy (anonymous):

?

OpenStudy (dayakar):

y = f(x) = x/2

OpenStudy (anonymous):

thats the first one

OpenStudy (dayakar):

B) f(x) =3x-10------(1) and f(x) = x/2-----(2) if x=8 plug in (1) and (2) compare (1) f(8) = 3*8-10 = 24-10 =14 can u do it same for (2) and compare